# Koopman Operators in Robot Learning

### v1

@@ VII Discussion and Conclusion
      Formal Accuracy Measures in General Hilbert and Banach Spaces: A significant limitation in Koopman-based modeling is the lack of formal accuracy measures applicable to general Hilbert and Banach spaces. Developing such measures is essential for optimization-based dictionary learning and ensuring the safety and reliability of applications in critical domains. These measures should be independent of the choice of basis for finite-dimensional subspaces, providing a standardized way to assess model accuracy and performance.
@@
      Identification of Classes of Lifting Functions: To ensure the effectiveness of Koopman-based control, it is crucial to identify classes of appropriate observable functions for lifting. While we have outlined different ways that have been proposed already, methods for higher-dimensional dynamical systems, or for cases where certain properties (like invariance) must be maintained, remain underdeveloped.
@@
      Sampling Rate Selection: One of the ongoing inquiries surrounding the application of Koopman-based approaches in robotic systems pertains to the selection of an optimal sampling rate. Particularly crucial for online learning, the sampling rate directly influences the quality of the derived Koopman operator. Additionally, it governs the error bounds associated with predictions. Selecting the rate for data collection and control signal execution will be vital to the performance of the robot.
@@
      Systems with Control Inputs: In forced systems, the linearity of observables does not automatically imply the linearity of the control inputs. For this reason, selected observables in existing works are often constrained to be linear functions of the input. However, this constraint can limit the quality of the resulting Koopman-based approximation since it may be too restrictive and fail to comprehensively capture the intricate interplay between input and state variables. A direct implication is that any convergence properties of the original control system may be altered, or even lost. In this regard, we believe there is much to be gained by exploiting the concept of the input-state separable form described earlier.
@@
      Uncertainty in Lifted Features: Lastly yet importantly, the assumption made in prior work regarding the zero-mean, Gaussian uncertainty model of the approximate Koopman operator should be revisited as well. Recent advancements have demonstrated that one can still operate and plan controllers that are robust to this specific choice of uncertainty model, as done in [144, 95]. Indeed, it is not clear if the nature of uncertainty in observables, especially with the addition of control input, can be accurately modeled as a normal Gaussian distribution. As an example, one might have a system where the states $x_{t}$ are well modeled as Gaussian (in this case the underlying dynamics might also be linear). It is not clear whether, under which conditions, Gaussian distributions pushed into the observable lifted space should form Gaussians. Thus, finding cases where lifting observables also preserves some statistical structure remains open.
    In conclusion, the application of Koopman operator theory in robotics represents a transformative shift in how we approach the modeling, control, and optimization of complex robotic systems in support of runtime learning. By leveraging the power of linear representation for inherently nonlinear dynamics, Koopman-based methods provide a robust and computationally efficient framework that addresses many of the limitations of traditional approaches. The theoretical advancements and practical implementations discussed in this review demonstrate the versatility and efficacy of Koopman operators across a wide array of robotic domains, including aerial, legged, wheeled, underwater, soft, and manipulator robots. As we continue to refine these techniques and explore new applications, the potential for Koopman operator theory to drive innovation in robotics is immense. Future research will undoubtedly focus on overcoming current challenges, such as incorporating constraints, improving real-time performance, and enhancing robustness to uncertainty. By doing so, we can fully harness the capabilities of Koopman theory to enable smarter, more adaptive, and more efficient robotic systems that can operate autonomously in complex, dynamic environments.

### v2

@@ VI Open Directions and Challenges
    Sampling Rate Selection: One of the ongoing inquiries surrounding the application of Koopman-based approaches in robotics pertains to the selection of an optimal sampling rate. Particularly crucial for online learning, the sampling rate directly influences the quality of the derived Koopman operator. Additionally, it governs the error bounds associated with predictions. Selecting the rate for data collection and control signal execution will be vital to the performance of the robot.
@@
    Extension to Fine/Dexterous Manipulation: The emergence of Koopman-based approaches for dexterous manipulation marks a promising expansion of this methodology. Contact-rich manipulation tasks, especially those involving fine-grained object interactions, non-prehensile control, and dexterous hand use, are becoming central challenges in robotics. Koopman operator theory shows strong potential as a solution, offering a global linear representation that can handle the discontinuities introduced by contact dynamics. Moreover, its data-driven nature enables the capture of subtle, object-specific behaviors that are difficult to model analytically, while still maintaining a structure conducive to efficient optimization. While this specific direction of research is still nascent, we believe it holds promise and merits more research.
@@
    Further Expansion on Soft Robotics: One fundamental limitation of Koopman-based linearization is the need to lift the system into a high-dimensional space, which can become computationally expensive or even intractable for continuum or soft-bodied systems. Since soft robots naturally exhibit high degrees of freedom, achieving effective linearization often requires dimensionality reduction techniques or simplified parametric representations. However, these approximations inevitably introduce biases and limit generalizability, particularly in contact-rich scenarios or tasks involving large deformations. How to choose the most appropriate description of a soft robot for compatibility with Koopman methods remains an open area of research.
@@
    Uncertainty in Lifted Features: Lastly yet importantly, the assumption made in prior work regarding the zero-mean, Gaussian uncertainty model of the approximate Koopman operator should be revisited as well. Recent advancements have demonstrated that one can still operate and plan controllers that are robust to this specific choice of uncertainty model, as done in [165, 82]. Indeed, it is not clear if the nature of uncertainty in observables, especially with the addition of control input, can be accurately modeled as a normal Gaussian distribution. As an example, one might have a system where the states $x_{t}$ are well modeled as Gaussian (in this case the underlying dynamics might also be linear). It is not clear whether, and under which conditions, Gaussian distributions pushed into the observable lifted space should form Gaussians. Thus, finding cases where lifting observables also preserves some statistical structure is a key direction of future research.
@@
    Study of the Underlying Optimization Problems: We have seen that the lifted system in the Koopman space is described by a linear (most often) or bilinear model, and that representing the (higher-dimensional) Koopman space observables with a (bi-)linear model does not conflict with the original (lower-dimensional) dynamics being nonlinear. While this property facilitates runtime adaptation, it may create challenges when solving the underlying optimization problems for control. Unless for special cases, the cost function can often depend on the lifted states. This can create challenges, e.g., in assessing the recursive feasibility of Koopman-NMPC methods. Although some challenges can be mitigated by enforcing a more careful implementation, how to design appropriate cost functions that can generalize across robotic systems remains an open problem.
VII Conclusion
In conclusion, the application of Koopman operator theory in robotics represents a transformative shift in how we approach the modeling, control, and optimization of complex robotic systems in support of runtime learning. By leveraging the power of a linear representation for inherently nonlinear dynamics, Koopman-based methods provide a robust and computationally efficient framework that addresses many of the limitations of traditional approaches. The theoretical advancements and practical implementations discussed in this review demonstrate the versatility and efficacy of Koopman operators across a wide array of robotic domains, including aerial, legged, wheeled, underwater, soft, and manipulator robots. We hope this survey can serve as an inviting gateway for newcomers, helping readers quickly build intuition for the theoretical underpinnings of the Koopman operator while also discovering its exciting applications across modern robotics. By demystifying the concepts and showcasing real-world use cases, we aim to spark curiosity, lower the barrier to entry, and empower more researchers and practitioners to explore, experiment, and advance this powerful framework. We believe Koopman-based methods hold tremendous potential, and we are excited to see how the robotics community will build on this foundation to drive future innovation.